# Spaces of subgroups of toral groups

### v1

@@ 16.E. Summary
If the module is $\mathbb{Z}[\omega]^{\prime}$ the extension is split and the finite full subgroups correspond to =$\mathcal{F}=\{n\in\mathbb{Z}\;|\;n\geq 1\}$ where $n$ gives the lattice $\Lambda^{S}$ generated by $n$, which corresponds to $H=\mathbb{T}[n]\rtimes\Sigma_{3}$.
If the module is $\mathbb{Z}[\omega]^{\prime\prime}$ and the extension is split, the finite full subgroups correspond to =$\mathcal{F}=\{n(2+\omega)\in\mathbb{Z}\;|\;n\geq 1\}$ where $n(2+\omega)$ gives the lattice $\Lambda^{S}$ generated by $n(2+\omega)$, which corresponds to $\mathbb{T}[n]\rtimes\Sigma_{3}$. In the non-split case $\mathcal{F}$ is replace by $\mathcal{F}_{0}$ of elements of $\mathcal{F}$ congruent to 0 mod 3.
@@ 17.A. Preamble
This example is the normalizer of the maximal torus inside a group of local type $G_{2}$ and dimension 14. In other words $\Lambda_{0}$ is the $G_{2}$-lattice and taken with the Lie theoretic Weyl group $W=D_{12}$ acting on the lattice. In concrete terms $D_{12}$ acts as the symmetry group of $\Lambda_{0}=\mathbb{Z}[\omega]$. A generating reflection is complex conjugation, and a generating rotation is by $62\pi/$ 6. We will use $1$ and $\omega$ as our basis elements.
@@ 17.C. Sections and conjugacy
$H^{1}(W;T^{2})\stackrel{{\scriptstyle a^{1}}}{{\longrightarrow}}H^{1}(W;T^{2}/S)\longrightarrow H^{2}(W;S)\stackrel{{\scriptstyle i^{S}_{*}}}{{\longrightarrow}}H^{2}(W;T^{2})\stackrel{{\scriptstyle a^{2}}}{{\longrightarrow}}H^{2}(W;T^{2}/S)\longrightarrow H^{3}(W;S)$

### v2

@@ 16.E. Summary
If the module is $\mathbb{Z}[\omega]^{\prime}$ the extension is split and the finite full subgroups correspond to $\mathcal{F}=\{n\in\mathbb{Z}\;|\;n\geq 1\}$ where $n$ gives the lattice $\Lambda^{S}$ generated by $n$, which corresponds to $H=\mathbb{T}[n]\rtimes\Sigma_{3}$.
If the module is $\mathbb{Z}[\omega]^{\prime\prime}$ and the extension is split, the finite full subgroups correspond to $\mathcal{F}=\{n(2+\omega)\in\mathbb{Z}\;|\;n\geq 1\}$ where $n(2+\omega)$ gives the lattice $\Lambda^{S}$ generated by $n(2+\omega)$, which corresponds to $\mathbb{T}[n]\rtimes\Sigma_{3}$. In the non-split case $\mathcal{F}$ is replace by $\mathcal{F}_{0}$ of elements of $\mathcal{F}$ congruent to 0 mod 3.
@@ 17.A. Preamble
This example is the normalizer of the maximal torus inside a group of local type $G_{2}$ and dimension 14. In other words $\Lambda_{0}$ is the $G_{2}$-lattice and taken with the Lie theoretic Weyl group $W=D_{12}$ acting on the lattice. In concrete terms $D_{12}$ acts as the symmetry group of $\Lambda_{0}=\mathbb{Z}[\omega]$. A generating reflection is complex conjugation, and a generating rotation is by $2\pi/6$. We will use $1$ and $\omega$ as our basis elements.
@@ 17.C. Sections and conjugacy
$H^{1}(W;T^{2})\stackrel{{\scriptstyle a^{1}}}{{\longrightarrow}}H^{1}(W;T^{2}/S)\longrightarrow H^{2}(W;S)\stackrel{{\scriptstyle i^{S}_{*}}}{{\longrightarrow}}H^{2}(W;T^{2})\stackrel{{\scriptstyle a^{2}}}{{\longrightarrow}}H^{2}(W;T^{2}/S)$